# Rise and fall of nonstabilizerness via random measurements

### v1

@@ I Introduction
Although much attention has been focused on the entanglement entropy [61, 62, 63, 64, 65, 66, 67, 68], increasing interest has turned to nonstabilizerness in monitored quantum circuits, exploring its dynamics under measurements. Nonstabilizerness has been studied in different settings such as Clifford+T circuits [57, 58], random quantum circuits [69, 70, 71] or monitored Gaussian fermions [72]. In Ref. [73], the authors investigated the decay of magic under sequences of local Cliffords and measurements, finding that the decay of magic can be characterized by two phases, depending on the rate of measurements. Further, Refs. [74, 75] studied the injection of magical noise channels within a random Clifford basis followed by measurements of most of the qubits, in analogy to quantum error correction under coherent noise, which is shown to remove magic below a critical noise rate. However, past works concentrated on the regime where the number of measurements is extensive with system size, while the role of individual measurements in the dynamics has not been well understood.
@@ II Nonstabilizerness
Let us consider $N-$qubit Pauli operators $P=\bigotimes_{i=1}^{N}\sigma^{\alpha}_{i}$ which are tensor products of single-qubit Pauli operators $\sigma^{x},\,\sigma^{y}$ and $\sigma^{z}$ and the single-qubit identity $\sigma^{0}=I_{1}$. We consider $\mathcal{P}_{N}$ as the set of all unsigned $N$-qubit Pauli operators $P\in\mathcal{P}_{N}$, that is the Pauli group modulo global phases {±$\{\pm 1,\pm i\}$, with cardinality $|\mathcal{P}_{N}|=4^{N}$. We denote $\mathcal{G}_{S}(|\psi\rangle)$ as the stabilizer group of a given state $|\psi\rangle$, i.e the set of commuting Pauli operators with
@@ III Protocol
Note that, due to the Clifford unitary randomizing the basis, we would get the equivalent dynamics if we choose any other qubit. We then measure the same qubit in the computational basis $\sigma_{1}^{z}$, record the measurement outcome {±$\in\{\pm 1\}$, and then apply the inverse Clifford unitary $U_{\text{C}}^{\dagger}$ on the full state. We repeat this process $t$ times, where we sample a random Clifford unitary $U_{\text{C}}$ at every step. Operationally, this process can be described by projection-valued measures (PVM)
@@
where $P^{\gamma}\equiv U_{C}^{\dagger}\sigma^{\gamma}_{1}U_{C}$ is a Pauli operator rotated into a random Clifford basis with $\gamma\in\{x,y,z\}$, and $\Pi_{b}$ is a generalized measure operator [4] satisfying $\sum_{b=\pm}\Pi_{b}^{{\dagger}}\Pi_{b}=I_{N}$. For each step, the probability of measuring the outcome ±$\pm 1$ in the state $|\psi\rangle$ is given by $p_{\pm}=\langle\psi|\Pi_{\pm}^{{\dagger}}\Pi_{\pm}|\psi\rangle$. After each measurement, we normalize the post-measurement state as $|\psi_{\pm}\rangle=\frac{1}{\sqrt{p_{\pm}}}\Pi_{\pm}|\psi\rangle$.
@@ IV.1 Analytical model
Note that the final state has unchanged nullity i.e. $\nu\rightarrow\nu$. Case $(iii)$ has $(4^{N-\nu}-2^{N-\nu})(4^{\nu}-1)$ possible Paulis. One can easily verify that $(i)$ to $(iii)$ cover all $P\in\mathcal{P}_{N}/\{I\}$ with in total $14^{N}-$ 1 Pauli operators.
@@
Our model now describes a Markov chain for the random variable $\nu$, with the update rule of the probability distribution $\rho(\nu)$ for each timestep $t$ given by
@@ IV.2 Numerical results
Figure 3: a) Change in SRE $\Delta\mathcal{M}_{2}^{\nu}$ at the $\nu$th measurement that yields a change in magic, averaged over $N_{\text{r}}\approx 1000$ trajectories, for different $N$ and for $\theta_{M}=0$. The dashed line is ¯$\overline{\mathcal{M}}_{\text{Haar}}^{\nu+1}-\overline{\mathcal{M}}_{\text{Haar}}^{\nu}$, where $\mathcal{M}_{\text{Haar}}^{N}$ is the SRE of Haar random state of $N$ qubits as defined in Eq. (26). b) Average SRE ¯$\overline{\mathcal{M}}_{2}$ over $N_{\text{r}}\approx 1000$ trajectories vs measurement step $t$ for three different initial states, $|\psi\rangle_{\text{Haar}}$, $|T\rangle^{\otimes N}$ and $|\psi\rangle_{\text{GUE}}$ for $\theta_{M}=0$. Inset: Average magic ¯$\overline{\mathcal{M}}_{2}$ for $|\psi\rangle_{\text{GUE}}$ with different axis range. c) Dots show numerical ¯$\overline{\mathcal{M}}_{2}$ against rescaled $\tau=A_{N}(t+1)$ for initial Haar-random states averaged over $N_{\text{r}}\approx 1000$ trajectories where $A_{N}\approx 2^{-N}$. Solid lines show ¯$\overline{\mathcal{M}}_{2}$ obtained using nullity from analytic prediction of the Markov chain in Eq. (20) inserted into SRE of Haar-random states of Eq. (26).
@@
We can now show that the model derived for the decay of nullity can also accurately capture that of the SRE for Haar-random states. Indeed, for such states, the Eq. (26) holds and maps the nullity $\nu\to\mathcal{M}_{2}$ at all times. In Fig. 3c, we show for Haar random states the average SRE ¯$\overline{\mathcal{M}}_{2}$ over $N_{r}\sim 1000$ trajectories as different points of different colors representing increasing $N$. The solid lines of the same color represent the mapping through Eq. (26) of the $\nu$ as predicted by the Markov chain rule in Eq. (20). Time is rescaled as $t\to A_{N}t$ (we again call $\tau=A_{N}(t+1)$) consistently with the scaling observed for stabilizer nullity as shown in Fig. 2, the different curves collapse onto a single universal one. The SRE of both $|T\rangle^{\otimes}$ and $|\psi\rangle_{\text{GUE}}$ states does not respect such functional relation, as further explained in Appendix D. Despite no analytical predictions on them can be made using this model, we can still use this relation as an effective model to phenomenologically describe SREs for different ensembles (in Appendix C).
@@ V.2 Numerical results
We now analyze the evolution of the average $\mathcal{M}_{2}$ for the same set of four different initial states in the presence of measurements in the magical basis ($\theta_{M}\neq 0$). Due to the non-monotonic behavior of SRE along individual trajectories, in the main text we primarily focus on the ensemble-averaged ¯$\overline{\mathcal{M}}_{2}$ and in particular, we consider the density $\overline{m}_{2}=\overline{\mathcal{M}}_{2}/N$. The behavior of individual trajectories for different $\theta_{M}$ and initial states is described in detail in Appendix A.
@@
The continuous growth of the steady state value $\overline{m}_{2}^{\text{SS}}$ with the angle parameter $\theta_{M}$ can be better visualized in Fig. 5b. We find that for ‘low-magic’ basis measurements ($\theta_{M}\ll 1$), the steady-state value increases quadratically with the angle, as shown in the same figure by the dashed fit lines. We find that the steady-state is in general not a Haar random state of the form Eq. (25), which is evidenced by the fact for any $\theta_{M}>0$ the nullity is close to $N-1$, while $\mathcal{M}_{2}$ can be relatively small especially for $\theta_{M}\ll 1$ [84]. However, the steady state of SREs appears to be close to the corresponding SRE of Haar random states of $N-2$ qubits, as shown in Appendix E.
@@
Figure 6: Average magic density of the steady state $\overline{m}_{2}^{\text{ss}}$ for $t\rightarrow\infty$ as a function of $\theta_{M}$ for different $N$ shown as dots. For $\theta_{\text{M}}\ll 1$, we find $\overline{m}_{2}^{\text{ss}}\propto\theta_{\text{M}}^{2}$ which is indicated as lines.
@@ VI Discussion
We then extended our analysis to single-qubit measurements in rotated bases, where measurements can also induce magic. Notably, rotated measurements can both generate and suppress magic, which converges to a steady-state for sufficiently long evolution [47]. We construct an analytical model for the stabilizer nullity $\nu$ in this generalized case, accurately capturing the stabilizer nullity of the ensemble steady state. Starting from an initial $\nu$, we observe that the average stabilizer nullity converges to about $N-1$, independent of $\theta_{M}$ and the initial state. Meanwhile, the steady-state value of the SRE $\mathcal{M}_{2}^{\text{ss}}$ depends on $\theta_{M}$, with a quadratic scaling $\mathcal{M}_{2}^{\text{ss}}\propto\theta_{M}^{2}$ in the small-angle limit. To support our analytical results, we performed extensive numerical simulations for various initial states, such as Haar-random states, unentangled product states (e.g., $T$-states) and short-time GUE-evolved states. Across all these initial conditions, we observed consistent behavior that confirms the validity of our conclusions and the generality of the observed phenomena.
@@ Appendix A Single trajectories in magical measurements
When measuring in the rotated bases, since the nullity evolution does not depend on the value of $\theta_{M}$, we can see that nullity will oscillate regardless of the angle around the asymptotic value $\nu^{\text{ss}}_{\text{analytic}}$ creating plateaus for $\nu=N-1\,,N-2$, (and with low probability $N-3$ and lower), which do not change with the angle. Two examples, one for initial Haar-random states and one for $|0\rangle^{\otimes N}$ states are shown in Fig. 7a,b. The dynamics does not depend on the value of $\theta_{M}$, and we show $\theta_{M}=0.0001$ to highlight the non-perturbative behavior of the stabilizer nullity.
@@
[…]
@@
Figure 9: Standard deviation $\sigma_{\mathcal{S}}$ of von Neumann entanglement entropy. a) $\overline{\sigma_{\mathcal{S}}}=\sigma_{\mathcal{S}}/\sqrt{N_{\text{r}}}$ vs $t+1$ for initial Haar-random states for $N=12$ qubits, $N_{r}=1000$, for different angle parameters fixed. b) $\overline{\sigma_{\mathcal{S}}}=\sigma_{\mathcal{S}}/\sqrt{N_{\text{r}}}$ vs $t+1$ for initial $|0\rangle^{\otimes N}$ states for $N=12$ qubits, $N_{r}=1000$.). c) Standard deviation $\overline{\sigma_{\mathcal{S}}}$ averaged over time $T$ as shown in Eq. (57). We plot against angle parameter $\theta_{M}$ for different qubit numbers $N$ for Haar-random initial states.
Next, we study traojectories of our circuit for the SRE. The behavior of the SRE is not very different from the nullity, except that the difference between two plateaus is not a fixed integer as shown in Fig. 3(a). Note that the measurement protocol with $\theta_{M}>0$ changes the single trajectories and therefore the average behavior. Also, it is worth noting that for every $\theta_{M}>0$ the nullity changes on timescales which depend linearly on $N$, while for the SRE we still need exponential timescales for $\theta_{M}<1$. In Fig. 5(a,b,c) we show how the SRE converges to the steady-state value for the different initial states, therefore in Fig. 10 we show how this average dynamics compares to single evolutions. For Haar-random states in Fig. 10(a,b,c) for increasing angle parameters, it is evident that the plateau structure the dynamics is conserved up to the moment the SRE reaches its angle-dependent steady-value. For $\theta_{M}=0.1$ in Fig. 10(b) we can clearly see that the moment the average approaches its steady value coincides with the moment the plateau structure is loose. For $\theta_{M}=1$ in Fig. 10(c), the SRE takes $\sim N$ measurements to reach $\mathcal{M}_{2}^{\text{ss}}$, which means there is no time to create plateaus.
@@ Appendix C Comparison numerical and analytic results
Figure 12: Comparison of the average stabilizer nullity $\overline{\nu}$ for different sizes with the approximate differential equation solution in Eq. (22) for different sizes $N$ as a function of measurement steps $t+1$.
@@
Figure 13: a) Comparison of the average stabilizer nullity $\overline{\nu}$ for different sizes with the approximate differential equation solution in Eq. (20) for different sizes $N$ as a function of measurement steps $t+1$, using $b_{f}=(1-y_{0,f})/(1+y_{0,f})$ and $A_{f}$ as fit parameters. b) Fit parameters $y_{0,f}$ and $A_{f}$ versus the number of qubits in the system, compared with analytical predictions. In particular, we can see that the predicted value of $y_{0,f}\sim 2^{N}$, which is $2^{\nu(0)}$. The predicted value of $A_{N}$ in Eq. (17) is displayed as yellow dashed line. As can be seen from the red dashed line, the fitted value is closer to being $2^{N+1}/(4^{N}-1)$. c) Average SRE ¯$\overline{\mathcal{M}}_{2}$ vs time for $\theta_{M}=0$ for $|\psi\rangle_{\text{Haar}}$, $|T\rangle^{\otimes N}$ and $|\psi\rangle_{\text{GUE}}$ for a system of $N=12$ qubits. For all states, dots are the numerical calculations, while for the lines we choose fitting ansatz inspired from our analytics. In particular, thet are the solution of the fitting procedure of Eq. (59) with free parameters $y_{0,f}$ and $A_{f}$.

### v2

@@ I Introduction
Although much attention has been focused on the entanglement entropy [61, 62, 63, 64, 65, 66, 67, 68], increasing interest has turned to nonstabilizerness in monitored quantum circuits, exploring its dynamics under measurements. Nonstabilizerness has been studied in different settings such as Clifford+T circuits [57, 58], random quantum circuits [69, 70, 71, 48] or monitored Gaussian fermions [72]. In Ref. [73], the authors investigated the decay of magic under sequences of local Cliffords and measurements, finding that the decay of magic can be characterized by two phases, depending on the rate of measurements. Further, Refs. [74, 75] studied the injection of magical noise channels within a random Clifford basis followed by measurements of most of the qubits, in analogy to quantum error correction under coherent noise, which is shown to remove magic below a critical noise rate. However, past works concentrated on the regime where the number of measurements is extensive with system size, while the role of individual measurements in the dynamics has not been well understood.
@@ II Nonstabilizerness
Let us consider NN-qubit Pauli operators $P=\bigotimes_{i=1}^{N}\sigma^{\alpha}_{i}$ which are tensor products of single-qubit Pauli operators $\sigma^{x},\,\sigma^{y}$ and $\sigma^{z}$ and the single-qubit identity $\sigma^{0}=I_{1}$. We consider $\mathcal{P}_{N}$ as the set of all unsigned $N$-qubit Pauli operators $P\in\mathcal{P}_{N}$, that is the Pauli group modulo global phases $\{\pm 1,\pm i\}$, with cardinality $|\mathcal{P}_{N}|=4^{N}$. We denote $\mathcal{G}_{S}(|\psi\rangle)$ as the stabilizer group of a given state $|\psi\rangle$, i.e the set of commuting Pauli operators with
@@ III Protocol
Note that, due to the Clifford unitary randomizing the basis, we would get the equivalent dynamics if we choose any other qubit. We then measure the same qubit in the computational basis $\sigma_{1}^{z}$, record the measurement outcome $\in\{\pm 1\}$. Note that here measurement and rotation are orthogonal, so nonstabilizerness can remain after measurement, in contrast to the case where rotation and measurement in the same basis as studied in Ref. [57].
Then, we apply the inverse Clifford unitary $U_{\text{C}}^{\dagger}$ on the full state. We repeat this process $t$ times, where we sample a random Clifford unitary $U_{\text{C}}$ at every step. Operationally, this process can be described by projection-valued measures (PVM)
@@
where $P^{\gamma}\equiv U_{C}^{\dagger}\sigma^{\gamma}_{1}U_{C}$ is a Pauli operator rotated into a random Clifford basis with $\gamma\in\{x,y,z\}$, and $\Pi_{b}$ is a generalized measure operator [4] satisfying $\sum_{b=\pm}\Pi_{b}^{{\dagger}}\Pi_{b}=I_{N}$. For each step, the probability of measuring the outcome $\pm 1$ in the state $|\psi\rangle$ is given by $p_{\pm}=\langle\psi|\Pi_{\pm}^{{\dagger}}\Pi_{\pm}|\psi\rangle$. After each measurement, we normalize the post-measurement state as $|\psi_{\pm}\rangle=\frac{1}{\sqrt{p_{\pm}}}\Pi_{\pm}|\psi\rangle$.
@@ IV.1 Analytical model
Note that the final state has unchanged nullity i.e. $\nu\rightarrow\nu$. Case $(iii)$ has $(4^{N-\nu}-2^{N-\nu})(4^{\nu}-1)$ possible Paulis. One can easily verify that $(i)$ to $(iii)$ cover all $P\in\mathcal{P}_{N}/\{I\}$ with in total $4^{N}-1$ Pauli operators.
@@
Our model now describes, in agreement with what derived in Ref. [48], a Markov chain for the random variable $\nu$, with the update rule of the probability distribution $\rho(\nu)$ for each timestep $t$ given by
@@ IV.2 Numerical results
[…]
@@
We can now show that the model derived for the decay of nullity can also accurately capture that of the SRE for Haar-random states. Indeed, for such states, the Eq. (26) holds and maps the nullity $\nu\to\mathcal{M}_{2}$ at all times. In Fig. 3c, we show for Haar random states the average SRE $\overline{\mathcal{M}}_{2}$ over $N_{r}\sim 1000$ trajectories as different points of different colors representing increasing $N$. The solid lines of the same color represent the mapping through Eq. (26) of the $\nu$ as predicted by the Markov chain rule in Eq. (20). Time is rescaled as $t\to A_{N}t$ (we again call $\tau=A_{N}(t+1)$) consistently with the scaling observed for stabilizer nullity as shown in Fig. 2, the different curves collapse onto a single universal one. The SRE of both $|T\rangle^{\otimes}$ and $|\psi\rangle_{\text{GUE}}$ states does not respect such functional relation, as further explained in Appendix D. Despite no analytical predictions on them can be made using this model, we can still use this relation as an effective model to phenomenologically describe SREs for different ensembles (in Appendix C).
@@ V.2 Numerical results
We now analyze the evolution of the average $\mathcal{M}_{2}$ for the same set of four different initial states in the presence of measurements in the magical basis ($\theta_{M}\neq 0$). Due to the non-monotonic behavior of SRE along individual trajectories, in the main text we primarily focus on the ensemble-averaged $\overline{\mathcal{M}}_{2}$ and in particular, we consider the density $\overline{m}_{2}=\overline{\mathcal{M}}_{2}/N$. The behavior of individual trajectories for different $\theta_{M}$ and initial states is described in detail in Appendix A.
@@
The continuous growth of the steady state value $\overline{m}_{2}^{\text{SS}}$ with the angle parameter $\theta_{M}$ can be better visualized in Fig. 5b. We find that for ‘low-magic’ basis measurements ($\theta_{M}\ll 1$), the steady-state value increases quadratically with the angle, as shown in the same figure by the dashed fit lines. We find that the steady-state is in general not a Haar random state of the form Eq. (25), which is evidenced by the fact for any $\theta_{M}>0$ the nullity is close to N1N-1, while $\mathcal{M}_{2}$ can be relatively small especially for $\theta_{M}\ll 1$ [84]. However, the steady state of SREs appears to be close to the corresponding SRE of Haar random states of N2N-2 qubits, as shown in Appendix E.
@@
Figure 6: Average magic density of the steady state $\overline{m}_{2}^{\text{ss}}$ for $t\rightarrow\infty$ as a function of $\theta_{M}$ for different NN shown as dots. For $\theta_{\text{M}}\ll 1$, we find $\overline{m}_{2}^{\text{ss}}\propto\theta_{\text{M}}^{2}$ which is indicated as lines.
@@ VI Discussion
We then extended our analysis to single-qubit measurements in rotated bases, where measurements can also induce magic. Notably, rotated measurements can both generate and suppress magic, which converges to a steady-state for sufficiently long evolution [47]. We construct an analytical model for the stabilizer nullity $\nu$ in this generalized case, accurately capturing the stabilizer nullity of the ensemble steady state. Starting from an initial $\nu$, we observe that the average stabilizer nullity converges to about N1N-1, independent of $\theta_{M}$ and the initial state. Meanwhile, the steady-state value of the SRE $\mathcal{M}_{2}^{\text{ss}}$ depends on $\theta_{M}$, with a quadratic scaling $\mathcal{M}_{2}^{\text{ss}}\propto\theta_{M}^{2}$ in the small-angle limit. Our findings for measurements in a rotated basis also provide a new lens through which to view the landscape of phases in monitored circuits. Of particular interest is the recently discussed phase of matter characterized by high, volume-law entanglement but low, area-law magic. [48]. Our model provides a tunable protocol to realize and stabilize such a state: by choosing $\theta_{M}\propto 1/\sqrt{N}$ our circuit naturally flows to a steady state with volume-law entanglement but constant non-zero magic. This provides a concrete and controllable setting to study the properties of this phase. To support our analytical results, we performed extensive numerical simulations for various initial states, such as Haar-random states, unentangled product states (e.g., $T$-states) and short-time GUE-evolved states. Across all these initial conditions, we observed consistent behavior that confirms the validity of our conclusions and the generality of the observed phenomena.
@@ Appendix A Single trajectories in magical measurements
When measuring in the rotated bases, since the nullity evolution does not depend on the value of $\theta_{M}$, we can see that nullity will oscillate regardless of the angle around the asymptotic value $\nu^{\text{ss}}_{\text{analytic}}$ creating plateaus for $\nu=N-1\,,N-2$, (and with low probability N3N-3 and lower), which do not change with the angle. Two examples, one for initial Haar-random states and one for $|0\rangle^{\otimes N}$ states are shown in Fig. 7a,b. The dynamics does not depend on the value of $\theta_{M}$, and we show $\theta_{M}=0.0001$ to highlight the non-perturbative behavior of the stabilizer nullity.
@@
$(\overline{\sigma_{\mathcal{S}}})_{T}=1/T\int_{t_{\text{min}}}^{t_{\text{max}}}\overline{\sigma_{\mathcal{S}}}dt\,$, (57)
@@
Figure 9: Standard deviation $\sigma_{\mathcal{S}}$ of von Neumann entanglement entropy. a) $\overline{\sigma_{\mathcal{S}}}=\sigma_{\mathcal{S}}/\sqrt{N_{\text{r}}}$ vs $t+1$ for initial Haar-random states for $N=12$ qubits, $N_{r}=1000$, for different angle parameters fixed. b) $\overline{\sigma_{\mathcal{S}}}=\sigma_{\mathcal{S}}/\sqrt{N_{\text{r}}}$ vs $t+1$ for initial $|0\rangle^{\otimes N}$ states for $N=12$ qubits, $N_{r}=1000$.). c) Standard deviation $\overline{\sigma_{\mathcal{S}}}$ averaged over time $T$ as shown in Eq. (57). We plot against angle parameter $\theta_{M}$ for different qubit numbers NN for Haar-random initial states.
Next, we study trajectories of our circuit for the SRE. The behavior of the SRE is not very different from the nullity, except that the difference between two plateaus is not a fixed integer as shown in Fig. 3(a). Note that the measurement protocol with $\theta_{M}>0$ changes the single trajectories and therefore the average behavior. Also, it is worth noting that for every $\theta_{M}>0$ the nullity changes on timescales which depend linearly on $N$, while for the SRE we still need exponential timescales for $\theta_{M}<1$. In Fig. 5(a,b,c) we show how the SRE converges to the steady-state value for the different initial states, therefore in Fig. 10 we show how this average dynamics compares to single evolutions. For Haar-random states in Fig. 10(a,b,c) for increasing angle parameters, it is evident that the plateau structure the dynamics is conserved up to the moment the SRE reaches its angle-dependent steady-value. For $\theta_{M}=0.1$ in Fig. 10(b) we can clearly see that the moment the average approaches its steady value coincides with the moment the plateau structure is loose. For $\theta_{M}=1$ in Fig. 10(c), the SRE takes $\sim N$ measurements to reach $\mathcal{M}_{2}^{\text{ss}}$, which means there is no time to create plateaus.
@@ Appendix C Comparison numerical and analytic results
Figure 12: Comparison of the average stabilizer nullity $\overline{\nu}$ for different sizes with the approximate differential equation solution in Eq. (22) for different sizes NN as a function of measurement steps $t+1$.
@@
Figure 13: a) Comparison of the average stabilizer nullity $\overline{\nu}$ for different sizes with the approximate differential equation solution in Eq. (20) for different sizes NN as a function of measurement steps $t+1$, using $b_{f}=(1-y_{0,f})/(1+y_{0,f})$ and $A_{f}$ as fit parameters. b) Fit parameters $y_{0,f}$ and $A_{f}$ versus the number of qubits in the system, compared with analytical predictions. In particular, we can see that the predicted value of $y_{0,f}\sim 2^{N}$, which is $2^{\nu(0)}$. The predicted value of $A_{N}$ in Eq. (17) is displayed as yellow dashed line. As can be seen from the red dashed line, the fitted value is closer to being $2^{N+1}/(4^{N}-1)$. c) Average SRE $\overline{\mathcal{M}}_{2}$ vs time for $\theta_{M}=0$ for $|\psi\rangle_{\text{Haar}}$, $|T\rangle^{\otimes N}$ and $|\psi\rangle_{\text{GUE}}$ for a system of $N=12$ qubits. For all states, dots are the numerical calculations, while for the lines we choose fitting ansatz inspired from our analytics. In particular, thet are the solution of the fitting procedure of Eq. (59) with free parameters $y_{0,f}$ and $A_{f}$.